### v1

@@ 16 Gaia benchmark stars, both dwarfs and giants
(blancocuaresma14, casamiquela25). Spectral lines of CCs are broader with respect to dwarfs and to red giants because of extended convective envelopes and pulsation,
which are parameterized by higher microturbulence () and
macroturbulence () values (da2023oxygen; luck81, luck85; bersier96).
@@
significant blending effects in CCs. Fig. 22 shows the difference between the atmospheric parameters derived in this work and the reference values for the Sun and the benchmark stars provided by blancocuaresma14.
@@
Definition of the continuum and RV optimization, we optimized the RV and the continuum level in each segment (, , );
Definition of the atmosphere model, the abundance scale and the NLTE grids. We used the grid of MARCS model atmospheres (gustafsson2008grid), that PySME-package defines as ”marcs2012”. We used solar abundances from asplund09 and the most updated versions of the NLTE grids 444https://zenodo.org/records/3888394 (amarsi2020galah). In order to have supersolar [N/Fe] and subsolar [C/Fe], typical of stars that have experienced mixing episodes during their evolution such as CCs, we used A(N)=8.23 and A(C)=8.2 for [Fe/H]=0.0;
@@
We focus on the elements for which NLTE grids are available in PySME, that are C, N, O, Na, Mg, Al, Si, S, K, Ca, Ti, Mn, Fe, Cu, Ba (amarsi2020galah, amarsi22, amarsi25; mallinson24; caliskan25). The line list highlighted in Table 3 is built from Gaia-ESO database according to the two quality parameters and (further details in Appendix C). We used a NARVAL solar spectrum and three UVES spectra of Red Giant Branch (RGB) stars of M67, in order to validate our line list on giant stars for which the chemical abundances are well established (occaso24). In PySME, we use the abundances of the specific element and the macroturbulence as free parameters. The final [X/Fe] value is the median value of the [X/Fe] values of each line. Table 4 includes the [X/Fe] values of the calibrating stars.
@@ 3.1 Atmospheric parameters
We estimated the atmospheric parameters in LTE once the NLTE grids were turned off, leaving the rest of the code unchanged. Fig. 2 shows the difference of the atmospheric parameters and of iron and titanium abundances derived under NLTE and LTE assumption, with histograms. Effective temperature, surface gravity and microturbulence velocity are the most affected by the departures, while Fe and Ti display minimal variations, indeed the differences are, on average smaller than -0.03 dex. Although, the difference between NLTE and LTE abundances is modest and the standard deviation is, on average, smaller than 0.08 dex, the individual differences can be of the order of dex. Individual differences are in several cases larger than random errors, thus suggesting that the NLTE approach gives narrower metallicity distribution functions (smaller standard deviations).
@@
Previous investigations addressing NLTE effects on the atmospheric parameters of FGK-type stars (see ruchti13; kovalev19) show that NLTE analyses generally yield higher , and values, particularly at low metallicities, as well as systematically higher microturbulent velocities of about for . Cepheids are supergiants and the is intrinsically much higher. The increase is supported by data plotted in Fig. 2, where is the only parameter that shows a statistically significant offset (), suggesting that microturbulence has a substantial impact on the determination of the other atmospheric parameters, which explains why we also find negative NLTE-LTE differences in , and . Further discussion and a comparison with previous studies is presented in Appendix D.
@@
To constrain possible differences with similar investigations we also compared our NLTE atmospheric parameters and abundances with those provided by da2023oxygen by using the same spectra (Fig. 3).
@@
The LTE atmospheric parameters provided by da2023oxygen were estimated by using pyMOOGi (adamow17), a Python version of the MOOG code (sneden73, sneden2002), atmosphere models from ATLAS (castelli2004new), and the ARES code (sousa07, sousa15) to measure the equivalent widths (EWs). In da2023oxygen, the authors used a large line list composed of 521 Fe I and Fe II lines based on different datasets and validated on calibrating Cepheids (proxauf2018new; da2022new). In contrast, our line list consists of 29 Fe I, Fe II, Ti I and Ti II, and was built in such a way as to be able to estimate atmospheric parameters via synthesis (see Sec 2.2 and Appendix C). Moreover, they derived by using the line depth ratio method (kovtyukh07; proxauf2018new), and , , and following the classical approach, i.e. consecutive iterations to reach ionization equilibrium for Fe I and Fe II lines and no trends in the Fe I abundances versus the EWs. We found positive trends in the difference of the atmospheric parameters, and indeed the slopes range from for the effective temperature (panel a) to for the surface gravity (panel b) and to for the microturbolence (panel c), but the mean differences are well within 1. The difference in iron abundance also shows a positive slope, which can be attributed both to differences in spectral analysis and to the fact that, as expected, the more metal-poor Cepheids appear even more metal-poor when analysed using a NLTE approach compared to LTE (bergemann2012non; hansen2013lte). However, the mean difference is quite small ().
@@
Furthermore, we compared our iron abundance with those from previous studies, specifically we do have 101 Cepheids in common with Trentin24B and 209 Cepheids in common with luck18. The top panel of Fig. 4 shows that the comparison with Trentin24B covers a broad range in iron abundance ( dex) and the difference does not show any trend, while the comparison with luck18 (bottom panel of the same figure) covers a narrower range in iron abundance ( dex), but the difference displays a clear positive trend. The mean differences are quite small 0.02 0.18 dex (Trentin24B) and -0.080.11 dex (luck18), but the individual differences, once again, reach 0.4 dex.
The reader interested in a more detailed discussion concerning the atmospheric parameters is referred to appendix C.
@@ 3.2 Iron radial gradient
The Galactocentric distances covered by our sample range from 4.6 kpc to 29.3 kpc, with four CCs located in the outskirt of the thin disk ( kpc).
To investigate the iron radial gradient we adopted three different analytical fitting functions (Fig. 5) by taking into account errors on both axes: , , and a bilinear function. Moreover, we adopted a Bayesian inference framework to estimate model parameters and their uncertainties. Specifically, given data and parameters , the posterior probability is:
where is the prior and the likelihood. We work with the log-posterior . We sampled the posterior with the affine-invariant Markov Chain Monte Carlo (MCMC) sampler implemented in the emcee library to obtain parameter point estimates (MAP) and credible intervals (the 68% credible region from the marginal posterior). In particular, we adopted mixture models (foremanmackey14), for which the per-data-point likelihood is given by a weighted sum of a foreground (signal) and a background (outlier) component. Writing for the mixing fraction (the prior probability that a datum belongs to the background component), the contribution of datum is:
and the total log-likelihood is . The parameter therefore controls the relative weight of the background/outlier model and robustifies the inference against points that are inconsistent with the foreground model. In Gaussian implementations the components are typically normal distributions with their own means and variances (optionally including an extra “jitter” term added in quadrature to account for underestimated measurement errors). Numerical evaluation of should use stable log-sum-exp arithmetic to avoid underflow when one component is much smaller than the other. Convergence of the MCMC chains was verified via autocorrelation-time estimates and visual inspection of trace plots.
@@
Data plotted in the top panel of Fig. 5 display quite clearly that the bilinear and the logarithmic fit reproduce quite well the radial gradient of CCs when moving from the innermost to the outermost disk regions. Note that the agreement for the bilinear fit is expected, since this fit has more degrees of freedom when compared with the linear and the logarithmic fit. However, the bilinear fit is also prone to possible systematics, since the edge of the two different intervals in Galactocentric distances is arbitrary. The maximized likelihood is obtained at a value of the ”knee” of (green dashed vertical line). In any case, the logarithmic fit is the most accurate analytical representation of the radial gradient, since the statistical parameters (MSE, reduced chi-squared, AIC score) attain their smallest values when compared with the other fits.
Fig. 6 shows the comparison between our linear iron radial gradient with those available in the literature, revealing a slope of -0.064 dex/kpc, which is steeper than the slopes provided by magrini23 by using open clusters and by luck18 by using CCs, and shallower than the slopes provided by Trentin24B based on CCs and by otto25 by using open clusters.
Fig. 7 displays the comparison of the current iron radial gradient with that of da2023oxygen in a plane. The slope of the current gradient is steeper (-1.53 vs. -0.91). The two radial profiles nearly overlap for kpc, but the current one becomes steeper in the outer disk. This difference is due to several reasons: the different spectral analysis method (including a different line list), the increased number of CCs at , which were not included in the da2023oxygen sample, and NLTE effects, since the outer disk Cepheids are the most metal-poor stars in the current sample.
To overcome possible problems in the choice of the adopted analytical function to describe the iron radial gradient, we decided to use the Gaussian Process Regression (GPR) to model the data without assuming a predefined functional form. The GPR is a Bayesian non-parametric method for modeling an unknown function in a regression setting, starting from a Gaussian process prior:
@@
We employed GPR by using Gpy555https://gpy.readthedocs.io/en/deploy/. After testing different kernel functions, we adopted GPy.kern.Matern32 according to the distribution dispersion and different sampling across the x-axis. Furthermore, we optimized the variance and the length-scale of the kernel function.
Data plotted in Fig. 8 display several interesting features worth being discussed in more detail. The GPR fits quite well the distribution of CCs as a function of the Galactocentric distance and this outcome applies to both the inner and the outer disk Cepheids. The GPR shows a particular trend in the data across the Solar circle, a ”shoulder” at 8 kpc. Moreover, GPR also shows a steady change in the slope for . This evidence indicates that the linear fit fails to take account for the radial gradient variations of the CCs in the outer disk, thus suggesting that the change is an intrinsic feature when moving into the outer disk. The radial distribution of open clusters younger than 400 Myr collected by otto25 agrees quite well with the radial distribution of Cepheids. This agreement is expected, since CCs are stellar tracers younger than 300 Myr. Unfortunately, the range in Galactocentric distances covered by this sample is too small when compared with CCs, but the agreement is very promising.
@@ 3.3 Radial gradients of chemical abundances
Distributions of the chemical abundance ratios are shown in Fig. 9, together with overlaid Gaussian fits (dashed red lines). The present results indicate Solar abundance ratios, within 1, for four -elements (O, Mg, Si, S) and for one light metal (Al). In contrast, [Na/Fe] ratio shows a clear overabundance of +0.300.12 dex, while [Cu/Fe] appears underabundant, with a mean value of dex. The Na overabundance is expected, since Cepheids are intermediate-mass stars and the low–mass tail experiences the first dredge-up along the red giant branch (see Fig. C13 in bono2024cepheids). A similar Na overabundance was also reported by Trentin24B, who found a mean value of dex, whereas their [Cu/Fe] measurements yielded a mean value of . otto25, analyzing open clusters who derived abundances for both main-sequence and giant stars, obtained mean values of dex and dex. In passing, an overabundance in [Na/Fe] was also found by genovali15.
Fig. 10 shows the radial abundance gradient of ten out of the eleven investigated chemical elements as a function of , with blue solid lines representing the GPR models. Similar to iron, a clear flattening of the gradients at kpc is obtained for all the elements, but for oxygen. Unfortunately, O abundances are only limited to CCs located within 14 kpc, since our Oxygen-determination is based on the triplet at 7770 which is only covered by FEROS, STELLA and ESPaDOnS spectra in our spectroscopic dataset.
@@
To further investigate the radial variation of the abundance gradients, Fig. 12 presents the different [X/Fe] ratios as a function of Galactocentric distance. The radial trends remain approximately constant for most of the investigated elements across the entire disk. The abundance ratio slopes are vanishing, suggesting that these elements follow the same radial behavior as iron. Three elements, however, deviate from this pattern: Al, Mn, and Cu. Mn and Cu display well-defined negative gradients (-0.0180.003, -0.0090.005 dex , respectively), while Al shows a positive gradient (+0.0120.003 dex ). The negative trend observed for Mn supports the behavior found for the iron-peak elements (Mn, Co, Ni) in OCs by otto25, whereas the nearly flat [X/Fe] ratios for O, Mg, Si, S, and Ca (all with slopes ¡0.01 dex ) further confirm the overall homogeneity of -element enrichment across the thin disk. genovali15 also reported nearly constant trends for Na, Al and Si. In contrast, they found mild evidence of positive slopes for Mg () and Ca (). However, their results were based on a smaller Cepheid sample, characterized by higher dispersion and a more limited spatial coverage.
@@
Fig. 13 shows for the investigated elements the abundance ratios as a function of (P) along with a linear fit. CCs obey to a well defined Period-Age relation (bono05, and references therein). The individual age steadily increases when moving from long- to short-period Cepheids. The abundance ratio of all the elements remain approximately constant across the full age range of the Cepheids in the sample. Only O, Na, Si, Ti, and Cu show a weak positive trend, 0.110.20, 0.100.03, 0.070.02, 0.090.02, 0.080.05, respectively. For oxygen, the higher dispersion and the shorter age coverage may affect the result. In other words, age-related effects are weak or negligible in the radial gradients and abundance ratios presented above. genovali15 found a slight negative trend for Ca (-0.15), in contrast with the null slope obtained in our study. In da2023oxygen, they found a slope of 0.080.02 in the [S/Fe] vs plane, while we obtained a weaker positive trend of 0.040.02. This suggests that the overall consistency of the abundances in our sample, which reduces the dispersion, allows for a more accurate assessment of whether a trend is present.
@@
In Fig. 14 the [X/Fe] vs [Fe/H] planes are shown. We emphasize the positive trend observed for Mn. Manganese is mainly produced in thermonuclear supernovae (Th-SNe), with yields that increase with metallicity (badenes08; kobayashi09), which explains the positive trend. Sodium, copper and aluminum are predominantly produced in core-collapse supernovae (CC-SNe); however, we observe a slight positive trend for Cu and a negative trend for Na and Al, reflecting their different metallicity dependences. kobayashi20 predict a constant or slightly negative trend for Na and Al, and a slight positive trend for Cu over the metallicity range we cover (), in agreement with our findings. Moreover, for Na and Al, the slight positive trend in [Na/Fe] and [Al/Fe] vs corresponds to a negative trend in the [Na/Fe] and [Al/Fe] vs [Fe/H] plane. Similarly, the negative trends of [Cu/Fe] and [Mn/Fe] with Galactocentric distance correspond to positive trends in the [Cu/Fe] and [Mn/Fe] vs [Fe/H] planes. For oxygen, magnesium, silicon, sulfur, calcium, and titanium, we expect a roughly constant trend in the metal-poor region of the sample and a decreasing trend at higher metallicities. This behavior arises because these elements are predominantly produced in CC-SNe, whereas metal-rich stars are increasingly affected by Th-SN enrichment, which reduces their [X/Fe] ratios. Since Si, S, Ca and Ti also receive contributions from Th-SNe, the negative trend in the metal-rich region is weak. In contrast, O and Mg are more affected and show a higher decreasing trend.
3.4 Radial gradients: LTE vs NLTE

### v2

@@ 16 Gaia benchmark stars, both dwarfs and giants
(Blanco-Cuaresma et al. 2014, Casamiquela et al. 2026). Spectral lines of CCs are broader with respect to dwarfs and to red giants because of extended convective envelopes and pulsation,
which are parameterised by higher microturbulence () and
macroturbulence () values (da Silva et al. 2023; Luck and Lambert 1981, 1985; Bersier and Burki 1996).
@@
significant blending effects in CCs. Fig. 22 shows the difference between the atmospheric parameters derived in this work and the reference values for the Sun and the benchmark stars provided by Blanco-Cuaresma et al. (2014), which take into account any NLTE corrections (Jofré et al. 2014).
@@
Definition of the continuum and RV optimisation, we optimised the RV and the continuum level in each segment (, , );
Definition of the atmosphere model, the abundance scale and the NLTE grids. We used the grid of MARCS model atmospheres (Gustafsson et al. 2008), that PySME-package defines as ’marcs2012’. We used solar abundances from Asplund et al. (2009) and the most updated versions of the NLTE grids 444https://zenodo.org/records/3888394 (Amarsi et al. 2020). In order to have supersolar [N/Fe] and subsolar [C/Fe], typical of stars that have experienced mixing episodes during their evolution such as CCs, we used A(N)=8.23 and A(C)=8.2 for [Fe/H]=0.0;
@@
A two-steps fitting procedure based on our selected line list is adopted to disentangle the contribution of each fitting parameters. In the first step, a simultaneous optimisation is performed to account for line broadening due to , which in CCs can reach substantial values (5-30 km or higher). This approach provides reliable estimates of and for the second step, while also yielding informed priors on , , and . By splitting the fitting into two lower-dimensional procedures, we reduced the computational cost while ensuring that the derived parameters were consistent with theoretical expectations for pulsating stars (see Appendix C).
We focus on the elements for which NLTE grids are available in PySME, that are C, N, O, Na, Mg, Al, Si, S, K, Ca, Ti, Mn, Fe, Cu, Ba (Amarsi et al. 2020, 2022, 2025; Mallinson et al. 2024; Caliskan et al. 2025). The line list highlighted in Table 12 is built from Gaia-ESO database according to the two quality parameters and (further details in Appendix C). We used a NARVAL solar spectrum and three UVES spectra of red clump stars of M67, in order to validate our line list on giant stars for which the chemical abundances are well established (Carbajo-Hijarrubia et al. 2024). In PySME, we use the abundances of the specific element and the macroturbulence as free parameters. The final [X/Fe] value is the median value of the [X/Fe] values of each line. Table 4 includes the [X/Fe] values of the calibrating stars.
@@ 3.1 Atmospheric parameters
We estimated the atmospheric parameters in LTE once the NLTE grids were turned off, leaving the rest of the code unchanged. Fig. 2 shows the difference of the atmospheric parameters and of iron and titanium abundances derived under NLTE and LTE assumption, with histograms. Effective temperature, surface gravity and microturbulence velocity are the most affected by the departures, while Fe and Ti abundances display minimal variations, indeed the differences are, on average smaller than dex. Although, the difference between NLTE and LTE abundances is modest and the standard deviation is, on average, smaller than 0.08 dex, the individual differences can be of the order of dex. Individual differences are in several cases larger than random errors, thus suggesting that the NLTE approach gives narrower metallicity distribution functions (smaller standard deviations).
@@
Previous investigations addressing NLTE effects on the atmospheric parameters of FGK-type stars (see Ruchti et al. 2013; Kovalev et al. 2019) show that NLTE analyses generally yield higher , and values, particularly at low metallicities, as well as systematically higher microturbulent velocities of about - for . Cepheids are supergiants and the is intrinsically much higher. The increase is supported by data plotted in Fig. 2, where is the only parameter that shows a statistically significant offset (), suggesting that microturbulence has a substantial impact on the determination of the other atmospheric parameters, which explains why we also find negative NLTE-LTE differences in , and . Further discussion and a comparison with previous studies is presented in Appendix D.
@@
To constrain possible differences with similar investigations we also compared our NLTE atmospheric parameters and abundances with those provided by dS23 by using the same spectra (Fig. 3).
@@
The LTE atmospheric parameters provided by dS23 were estimated by using pyMOOGi (Adamow 2017), a Python version of the MOOG code (Sneden 1973, 2002), atmosphere models from ATLAS (Castelli and Kurucz 2004), and the ARES code (Sousa et al. 2007, 2015) to measure the equivalent widths (EWs). In dS23, the authors used a large line list composed of 521 Fe I and Fe II lines based on different datasets and validated on calibrating Cepheids (Proxauf et al. 2018; da Silva et al. 2022). In contrast, our line list consists of 36 Fe I, Fe II, Ti I and Ti II, and was built in such a way as to be able to estimate atmospheric parameters via synthesis (see Sec 2.2 and Appendix C). Moreover, they derived by using the line depth ratio method (Kovtyukh 2007; Proxauf et al. 2018), and , , and following the classical approach, i.e. consecutive iterations to reach ionization equilibrium for Fe I and Fe II lines and no trends in the Fe I abundances versus the EWs. We found positive trends in the difference of the atmospheric parameters, and indeed the slopes range from for the effective temperature (panel a) to for the surface gravity (panel b) and to for the microturbolence (panel c), but the mean differences are well within 1. The difference in iron abundance also shows a positive slope, which can be attributed both to differences in spectral analysis and to the fact that, as expected, the more metal-poor Cepheids appear even more metal-poor when analysed using a NLTE approach compared to LTE (Bergemann et al. 2012; Hansen et al. 2013). However, the mean difference is quite small ().
@@
Furthermore, we compared our iron abundance with those from previous studies, specifically we do have 101 Cepheids in common with Trentin et al. (2024a) and 209 Cepheids in common with Luck (2018). The top panel of Fig. 4 shows that the comparison with Trentin et al. (2024a) covers a broad range in iron abundance ( dex) and the difference does not show any trend, while the comparison with Luck (2018) (bottom panel of the same figure) covers a narrower range in iron abundance ( dex), but the difference displays a clear positive trend. The mean differences are quite small dex (Trentin et al. 2024a) and dex (Luck 2018), but the individual differences, once again, reach 0.4 dex.
The trends and offsets observed in comparison with the literature (Figs. 3 and 4) reflect the methodological nature of the discrepancies, which arise primarily from the adoption of different fitting strategies and line lists, and secondarily differences in the underlying physical assumption. This conclusion is supported by the tests presented in Appendix C and Appendix D, where a more detailed discussion of the atmospheric parameters is provided.
@@ 3.2 Iron radial gradient
The Galactocentric distances covered by our sample range from 4.6 kpc to 29.3 kpc, with four CCs located in the outskirt of the thin disk ( kpc). To investigate the iron radial gradient we adopted three different analytical fitting functions (Fig. 5) by taking into account errors on both axes: , , and a bilinear function.
Moreover, we adopted a Bayesian inference framework to estimate model parameters and their uncertainties. Specifically, given data and parameters , the posterior probability is:
where is the prior and the likelihood. We work with the log-posterior . We sampled the posterior with the affine-invariant Markov Chain Monte Carlo (MCMC) sampler implemented in the emcee library to obtain parameter point estimates (MAP) and credible intervals (the 68% credible region from the marginal posterior). In particular, we adopted mixture models (Foreman-Mackey 2014), for which the per-data-point likelihood is given by a weighted sum of a foreground (signal) and a background (outlier) component. Writing for the mixing fraction (the prior probability that a datum belongs to the background component), the contribution of datum is:
and the total log-likelihood is . The parameter therefore controls the relative weight of the background/outlier model and robustifies the inference against points that are inconsistent with the foreground model. In Gaussian implementations the components are typically normal distributions with their own means and variances (optionally including an extra ’jitter’ term added in quadrature to account for underestimated measurement errors). Numerical evaluation of should use stable log-sum-exp arithmetic to avoid underflow when one component is much smaller than the other. Convergence of the MCMC chains was verified via autocorrelation-time estimates and visual inspection of trace plots.
@@
Data plotted in the top panel of Fig. 5 display quite clearly that the bilinear and the logarithmic fit reproduce quite well the radial gradient of CCs when moving from the innermost to the outermost disk regions. Note that the agreement for the bilinear fit is expected, since this fit has more degrees of freedom when compared with the linear and the logarithmic fit. However, the bilinear fit is also prone to possible systematics, since the edge of the two different intervals in Galactocentric distances is arbitrary. The maximized likelihood is obtained at a value of the ’knee’ of (green dashed vertical line). In any case, the logarithmic fit is the most accurate analytical representation of the radial gradient, since the statistical parameters (MSE, reduced chi-squared, AIC score) attain their smallest values when compared with the other fits.
Fig. 6 shows the comparison between our linear iron radial gradient with those available in the literature, revealing a slope of dex/kpc, which is steeper than the slopes provided by Magrini et al. (2023, ) by using open clusters and by Luck (2018, ) by using CCs, and shallower than the slopes provided by Trentin et al. (2024a, ) based on CCs and by Otto et al. (2025, ) by using open clusters.
Fig. 7 displays the comparison of the current iron radial gradient with that of dS23 in a plane. The slope of the current gradient is steeper ( vs ). The two radial profiles nearly overlap for kpc, but the current one becomes steeper in the outer disk. This difference is due to several reasons: the different spectral analysis method (including a different line list), the increased number of CCs at , which were not included in the dS23 sample, and NLTE effects, since the outer disk Cepheids are the most metal-poor stars in the current sample.
To overcome possible problems in the choice of the adopted analytical function to describe the iron radial gradient, we decided to use the Gaussian Process Regression (GPR) to model the data without assuming a predefined functional form. The GPR is a Bayesian non-parametric method for modelling an unknown function in a regression setting, starting from a Gaussian process prior:
@@
We employed GPR by using Gpy555https://gpy.readthedocs.io/en/deploy/. After testing different kernel functions, we adopted GPy.kern.Matern32 according to the distribution dispersion and different sampling across the x-axis. Furthermore, we optimised the variance and the length-scale of the kernel function.
Data plotted in Fig. 8 display several interesting features worth being discussed in more detail. The GPR fits quite well the distribution of CCs as a function of the Galactocentric distance and this outcome applies to both the inner and the outer disk Cepheids. The GPR shows a particular trend in the data across the Solar circle, a ’shoulder’ at 8 kpc. Moreover, GPR also shows a steady change in the slope for 14-16 kpc. This evidence indicates that the linear fit fails to take account the radial gradient variations of the CCs in the outer disk, thus suggesting that the change is an intrinsic feature when moving into the outer disk. The radial distribution of open clusters younger than 400 Myr collected by Otto et al. (2025) agrees quite well with the radial distribution of Cepheids. This agreement is expected, since CCs are stellar tracers younger than 300 Myr. Unfortunately, the range in Galactocentric distances covered by this sample is too small when compared with CCs, but the agreement is very promising.
@@ 3.3 Radial gradients of chemical abundances
Distributions of the chemical abundance ratios are shown in Fig. 9, together with overlaid Gaussian fits (dashed red lines). The present results indicate Solar abundance ratios, within 1, for four -elements (O, Mg, Si, S) and for one light metal (Al). In contrast, [Na/Fe] ratio shows a clear overabundance of +0.300.12 dex, while [Cu/Fe] appears underabundant, with a mean value of dex. The Na overabundance is expected, since Cepheids are intermediate-mass stars and the low–mass tail experiences the first dredge-up along the red giant branch (see Fig. C13 in Bono et al. 2024). A similar Na overabundance was also reported by Trentin et al. (2024a), who found a mean value of dex, whereas their [Cu/Fe] measurements yielded a mean value of . Otto et al. (2025), analyzing open clusters who derived abundances for both main-sequence and giant stars, obtained mean values of dex and dex. In passing, an overabundance in [Na/Fe] was also found by Genovali et al. (2015).
Figure 10 shows the radial abundance gradient of ten out of the eleven investigated chemical elements as a function of , with blue solid lines representing the GPR models. Similar to iron, a clear flattening of the gradients at 14-16 kpc is obtained for all the elements, but for oxygen. Unfortunately, O abundances are only limited to CCs located within 14 kpc, since our oxygen-determination is based on the triplet at 7770 which is only covered by FEROS, STELLA and ESPaDOnS spectra in our spectroscopic dataset.
@@
To further investigate the radial variation of the abundance gradients, Fig. 12 presents the different [X/Fe] ratios as a function of Galactocentric distance. The radial trends remain approximately constant for most of the investigated elements across the entire disk. The abundance ratio slopes are vanishing, suggesting that these elements follow the same radial behaviour as iron. Three elements, however, deviate from this pattern: Al, Mn, and Cu. Mn and Cu display well-defined negative gradients (, dex , respectively), while Al shows a positive gradient ( dex ). The negative trend observed for Mn supports the behaviour found for the iron-peak elements (Mn, Co, Ni) in OCs by Otto et al. (2025), whereas the nearly flat [X/Fe] ratios for O, Mg, Si, S, and Ca (all with slopes ¡0.01 dex ) further confirm the overall homogeneity of -element enrichment across the thin disk. Genovali et al. (2015) also reported nearly constant trends for Na, Al and Si. In contrast, they found mild evidence of positive slopes for Mg () and Ca (). However, their results were based on a smaller Cepheid sample, characterized by higher dispersion and a more limited spatial coverage.
@@
Figure 13 shows for the investigated elements the abundance ratios as a function of (P) along with a linear fit. CCs obey to a well defined Period-Age relation (Bono et al. 2005, and references therein). The individual age steadily increases when moving from long- to short-period Cepheids. The abundance ratio of all the elements remain approximately constant across the full age range of the Cepheids in the sample. Only O, Na, Si, Ti, and Cu show a weak positive trend, 0.110.20, 0.100.03, 0.070.02, 0.090.02, 0.080.05, respectively. For oxygen, the higher dispersion and the shorter age coverage may affect the result. In other words, age-related effects are weak or negligible in the radial gradients and abundance ratios presented above. Genovali et al. (2015) found a slight negative trend for Ca (-0.15), in contrast with the null slope obtained in our study. In dS23, they found a slope of 0.080.02 in the [S/Fe] versus plane, while we obtained a weaker positive trend of 0.040.02. This suggests that the overall consistency of the abundances in our sample, which reduces the dispersion, allows for a more accurate assessment of whether a trend is present.
@@
In Fig. 14 the [X/Fe] vs [Fe/H] planes are shown. We emphasize the positive trend observed for Mn. Manganese is mainly produced in thermonuclear supernovae (Th-SNe), with yields that increase with metallicity (Badenes et al. 2008; Kobayashi and Nomoto 2009), which explains the positive trend. Sodium, copper and aluminum are predominantly produced in core-collapse supernovae (CC-SNe); however, we observe a slight positive trend for Cu and a negative trend for Na and Al, reflecting their different metallicity dependences. Kobayashi et al. (2020) predict a constant or slightly negative trend for Na and Al, and a slight positive trend for Cu over the metallicity range we cover (), in agreement with our findings. Moreover, for Na and Al, the slight positive trend in [Na/Fe] and [Al/Fe] versus corresponds to a negative trend in the [Na/Fe] and [Al/Fe] versus [Fe/H] plane. Similarly, the negative trends of [Cu/Fe] and [Mn/Fe] with Galactocentric distance correspond to positive trends in the [Cu/Fe] and [Mn/Fe] versus [Fe/H] planes. For oxygen, magnesium, silicon, sulfur, calcium, and titanium, we expect a roughly constant trend in the metal-poor region of the sample and a decreasing trend at higher metallicities. This behaviour arises because these elements are predominantly produced in CC-SNe, whereas metal-rich stars are increasingly affected by Th-SN enrichment, which reduces their [X/Fe] ratios. Since Si, S, Ca and Ti also receive contributions from Th-SNe, the negative trend in the metal-rich region is weak. In contrast, O and Mg are more affected and show a higher decreasing trend.
3.4 Radial gradients: LTE versus NLTE